OpenStudy (anonymous):

Use the distributive property to rewrite the expression –2(6 – x) without parentheses. A. –8 + 2x B. 2x – 12 C. –12 – 2x D. –10x

jigglypuff314 (jigglypuff314):

distributive property is like a(b+c) = a*b + a*c so -2(6 - x) = (-2)*(6) + (-2)*(-x)